OpenStudy (anonymous):

find the 6th term of the geometric sequence where t1=2 and r=5

OpenStudy (amistre64):

a geometric sequence is just multiplying each term by r to get the next term start at t1 t1 = 2 t2 = t1 * 5 t3 = t2 * 5 t4 = t3 * 5 and so on